# Operator Deep Smoothing for Implied Volatility

### v1

@@ Abstract
We devise a novel method for implied volatility smoothing based on neural operators. The goal of implied volatility smoothing is to construct a smooth surface that links the collection of prices observed at a specific instant on a given option market. Such price data arises highly dynamically in ever-changing spatial configurations, which poses a major limitation to foundational machine learning approaches using classical neural networks. While large models in language and image processing deliver breakthrough results on vast corpora of raw data, in financial engineering the generalization from big historical datasets has been hindered by the need for considerable data pre-processing. In particular, implied volatility smoothing has remained an instance-by-instance, hands-on process both for neural network-based and traditional parametric strategies. Our general operator deep smoothing approach, instead, directly maps observed data to smoothed surfaces. We adapt the graph neural operator architecture to do so with high accuracy on ten years of raw intraday S&P 500 options data, using a single set of weights. The trained operator adheres to critical no-arbitrage constraints and is robust with respect to subsampling of inputs (occurring in practice in the context of outlier removal). We provide extensive historical benchmarks and showcase the generalization capability of our approach in a comparison with SVI, an industry standard parametrization for implied volatility. The operator deep smoothing approach thus opens up the use of neural networks on large historical datasets in financial engineering.
@@ 1 Introduction
Options trading experienced phenomenal growth in recent years. In its 2023 trading volume report [1], the CBOE announced the fourth consecutive year of record-breaking volumes on its options exchanges, citing a record-breaking number of transactions for European options on the S&P 500 index. European options are financial derivative contracts that give their holder the right, but not the obligation, to either buy or sell an underlying asset at a predetermined price (the strike) at a predetermined time (the expiry). An option specifying the right to buy (respectively to sell) is called a Call (respectively Put) option. Options are traded on a wide range of underlyings, including stocks, indices, currencies and commodities, and can be used to hedge against or speculate on the price movements of the underlying asset.
@@
A key concept in options trading is the so-called implied volatility, which transforms the nominal price of an option into a conceptually and numerically convenient metric. The implied volatility surface is the collection of implied volatilities as observed at a specific point in time, visualized in three-dimensional space as a surface over the (strike, expiry)-domain. It provides an intuitive representation of the state of the options market and is crucial for hedging and risk management. The extraction of a smooth surface from quoted option prices is called implied volatility smoothing and remains one of the key challenges in options trading.
@@
Conventionally, implied volatility smoothing relies on parametric surfaces whose parameters are optimized based on the distance to observed prices while adhering to absence-of-arbitrage conditions, which ensure the consistency of prices extrapolated from the smoothed surface. The development of such ad-hoc models for implied volatility traces back to SVI [10], which models implied volatility slice-wise for each maturity and successfully captures its key features on Equity indices. A continuous interpolation scheme for SVI slices was provided in [11], yielding a full surface. Nowadays, sophisticated market makers employ custom parametrizations, which can be considered proprietary trading secrets and reduce SVI to a benchmark role.
@@
Regardless of the particular parametric surface model used, these approaches boil down to the continued execution of a numerical optimization routine: A smoothed surface expires as soon as quotes are updated (whenever markets move), necessitating the re-calibration of parameters. Success and duration of this routine is sensitive to initial conditions, search heuristics, and termination criteria, which exposes practitioners to considerable uncertainties during trading hours (or online). In response, we introduce a novel operator deep smoothing approach, replacing the instance-by-instance optimization with a single evaluation of a neural network. This greatly simplifies online calibration, at the upfront cost of training the network offline from historical data (in the spirit of [16, 17, 25]). Our unique use of neural operators [20] is fundamentally directed by the observation that the raw inputs for volatility smoothing (the collections of observed volatilities) vary in size and spatial arrangement: Options expire, new maturities and strikes become available, and the coordinates of existing options evolve continuously in the domain of the implied volatility surface (Figure 1(b)). This setting excludes classical neural networks – which required fixed-size inputs – from direct application. Neural operators, instead, conceptualize observed data as point-wise discretizations of latent functions in implicit infinite-dimensional function spaces and are well suited for the task. As the aim of volatility smoothing is the construction of a smooth surface that links the collection of market-observed option quotes, their use is highly motivated.
@@
We introduce operator deep smoothing, a general approach for discretization-invariant data interpolation based on neural operators, and apply it to implied volatility smoothing. Our technique transcends traditional parametric smoothing and directly maps quoted volatilities to smoothed surfaces. Standard methods are limited to certain option markets (e.g. FX markets where options spread out on fixed rectilinear grids [5]) or require substantial data pre-processing. Instead, our technique adapts the graph neural operator (GNO) architecture [3] to consistently smooth input data of variable number and spatial arrangement. While neural operators have been successfully used in Physics to numerically solve partial differential equations, our application is the first in Finance and highlights the values of their discretization-invariance properties, so far rather under-explored. We test our method on ten years of intraday S&P 500 options data, accounting for more than 60 million volatility datapoints. Our method improves on the errors reported in [2], where much shorter periods of data were considered. In contrast, here we employ nine years of training data, and obtain double performance compared to the industry benchmark SVI on the final year of data. We summarize the impact of our contributions as follows:
  • Neural Operators in Financial Engineering – Our GNO approach consistently maps quoted volatilities to smooth surfaces and is well suited to handle the highly dynamic nature of options data, removing the need for instance-by-instance processing of incoming data. Surfaces (or higher-dimensional equivalents) similar to the implied volatility surface are ubiquitous in quantitative finance, and we expect our technique to be transferable and to streamline and robustify algorithms’ data pipelines.
@@
  • Neural Operators for Discretization-Invariant Interpolation – Our operator deep smoothing approach leverages (a tweaked version of) GNOs for their abilities to interpolate irregular mesh geometries. This constitutes the first application of neural operators for interpolation/extrapolation tasks and paves the way to future research on the versatility of the discretization-invariance of neural operators in industrial applications characterized by dynamic and spatially irregular data.
Broader Impact
Financial markets are centerpieces of modern economies. Operator deep smoothing, an algorithm to achieve robust and realistic generalization from market data, can help financial intermediaries monitor and control financial risks better, thereby contributing to improved market efficiency. The beneficiaries include entities like pension funds, who use options to safeguard payment plans, and mutual funds, seeking to manage portfolio risks efficiently. Ultimately, operator deep smoothing may be useful for all market participants who own options in their portfolios. This includes the general public, whose trading participation has been increasing substantially [8]. Operator deep smoothing, providing "cheap" implied volatility smoothing with a single evaluation of a neural network, can thus help improve the quality of broadly accessible investment tools.
@@
The aforementioned SVI was developed for internal use at Merrill Lynch in 1999 and later advocated in [10]. Its extension to surface-based SSVI in [11] has been eagerly adopted by practitioners, which have since contributed to its robust calibration and generalizations [7, 15, 12]. It was augmented in [2] by a multiplicative neural network corrector, based on guided network training by means of no-arbitrage soft constraints from [34]. The absence-of-arbitrage conditions – providing safeguards for option pricing – for implied volatility surfaces were formulated in [31], and we provide an equivalent formulation, based on [9, 26], for practical purposes. In [6] static arbitrage constraints were used to perform option calibration (with an additional regularization technique), which can be considered to be instance-by-instance smoothing of nominal price data. In [5] a classical VAE (variational autoencoder) was applied to implied volatility smoothing on FX markets, where strikes of quoted options are tied to a fixed grid of deltas. [^1] Delta is the derivative of the price with respect to the underlying asset and is standard in FX strike quoting. This specificity of FX markets allows the use of a conventional feedforward neural network based decoder. Recent option calibration approaches based on neural networks have been proposed in [4, 16, 17, 33].
A comprehensive account on neural operators is given in [20], unifying previous research on different neural operator architectures and techniques [3, 22]. Subsequent developments investigating the expressivity of these architectures as well as their generalizations include [13, 18, 21, 23, 24, 32].
@@
We review financial concepts and the challenges of implied volatility smoothing in Section 2. In Section 3, we provide a review of neural operators (Section 3.1) and introduce (Section 3.2) our operator deep smoothing approach for general interpolation tasks. In Section 4, we perform experiments for implied volatility smoothing of S&P 500 options data. Finally, Section 5 gathers limitations as well as outlooks regarding the use of neural operators for interpolation purposes.
@@
We make code for the paper available at the location https://github.com/rwicl/operator-deep-smoothing-for-implied-volatility. In particular, the code repository contains a general PyTorch [30] implementation of the graph neural operator architecture for operator deep smoothing.
@@ 2 Background: Implied Volatility
We consider a market of European options written on an underlying asset, which we observe at a given instant $T_{0}$ and denote the time-$T$ forward price of the underlying asset by $F_{T_{0},T}$.
@@
This model also provides a closed-form expression for the sensitivity of the option price with respect to its volatility parameter. Termed Vega and required for our experiments in Section 4.1, it reads
@@
While not able (any longer) to fit market data, the mathematical tractability of the Black-Scholes model gave rise to the important concept of implied volatility: Given a Call option with price $C(T,K)$, its implied volatility $v(\tau,k)$ solves $C(T,K)=F_{T_{0},T}\operatorname{BS}(\tau,k,v(\tau,k))$. By using time-to-expiry/log-moneyness coordinates, the implied volatility provides a universal way to consistently compare the relative expensiveness of options of across different strikes, maturities, underlyings, and interest rate environments. Its characteristic shape helps traders make intuitive sense of the instantaneous state of option markets relative to a (flat) Black-Scholes model baseline.
@@
This refers to fitting a smooth surface $\hat{v}\colon(0,\infty)\times\mathbb{R}\to(0,\infty)$ to a collection $\mathbf{v}=\{v(\tau_{l},k_{l})\}_{l=1}^{p}$ of observed implied volatilities. Naive strategies such as cubic interpolation, are ill-fated as they yield arbitrage, namely cost-less trading strategies generating a guaranteed profit. In option markets, an arbitrage is called static when set up solely from fixed positions in options and a dynamically (but a finite number of times) readjusted position in the underlying. Beyond simple interpolations, practitioners have devised ad-hoc parametrizations for implied volatility, in particular the aforementioned SVI, which are not expected to perfectly match all reference prices. Instead, the model parameters are optimized with respect to an objective function that measures market price discrepancy and includes penalization terms ruling out static arbitrage. These penalization terms are commonly formulated on the basis of the following theorem, which summarizes the shape constraints of the implied volatility surface [11, 26, 31].
@@
Condition 2.1(i) is equivalent to prices increasing in maturity (uncertain increases as time passes), while Condition 2.1(ii) arises when computing the implied probability density $f_{\tau}$ of the underlying:
@@
Since a density needs to be non-negative, Identity (4) explains why Condition 2.1(ii) above is required.
(a) Relative trading volume per quoted interval, averaged over S&P 500 dataset 2012-2021. Left: Rectangular domain w.r.t.  time-to-expiry/log-moneyness (99.9% of trading volume with maturities <1 year). Right: Rectangular domain w.r.t. transformed coordinates (96.6% of trading volume while more evenly populated); boundary delineated in Left.
@@ 3.1 Background: Neural Operators
An input-output relationship of data $\mathbf{a}\mapsto\mathbf{u}$ is then "really" described by an operator $F\colon\mathcal{A}\to\mathcal{U}$ between function spaces $\mathcal{A}$ and $\mathcal{U}$. Neural operators are abstract neural network architectures $F^{\theta}\colon\mathcal{A}\to\mathcal{U}$, with implementations that integrate (5) consistently across the variable discretization $\pi$.
@@
We review the core concepts of neural operators from [20]. Let $D$ be a bounded domain in $\mathbb{R}^{d}$ and $\mathcal{A}$ and $\mathcal{U}$ be Banach spaces of functions mapping from $D$ to $\mathbb{R}^{c_{\text{in}}}$ and $\mathbb{R}^{c_{\text{out}}}$, respectively. Neural operators are finitely parametrized mappings $F^{\theta}\colon\mathcal{A}\to\mathcal{U}$ with universality for continuous target operators and with discretization-invariant implementations $\tilde{F}^{\theta}$. In the space $C(\mathcal{A},\mathcal{U})$ topologized by uniform convergence on compacts, the architecture $F^{\theta}$ is called universal if $\{F^{\theta}\}_{\theta\in\Theta}$ is dense in $C(\mathcal{A},\mathcal{U})$, with $\Theta$ the parameter set. An implementation of $F^{\theta}$ is an algorithm $\tilde{F}^{\theta}$ which accepts observed data $\mathbf{a}=a|_{\pi}$ and outputs a function $u\in\mathcal{U}$ and is such that $\tilde{F}^{\theta}_{\pi}(\makebox[4.30554pt]{{$\cdot$}})=\tilde{F}^{\theta}(\makebox[4.30554pt]{{$\cdot$}}|_{\pi})\in C(\mathcal{A},\mathcal{U})$. Now, $\tilde{F}^{\theta}$ is called discretization-invariant if $\lim_{n\uparrow\infty}\tilde{F}^{\theta}_{\pi^{(n)}}=F^{\theta}$ in $C(\mathcal{A},\mathcal{U})$, given a discrete refinement [^3] A discrete refinement of $D$ is a nested sequence $(\pi^{(n)})_{n\in\mathbb{N}}$ of discretizations of $D$ for which for every $\varepsilon>0$, there exists $N\in\mathbb{N}$ such that $\{\mathbb{B}_{\mathbb{R}^{d}}(x,\varepsilon)\colon x\in\pi^{(N)}\}$ covers $D$. of $D$.

### v2

@@ Abstract
We devise a novel method for nowcasting implied volatility based on neural operators. Better known as implied volatility smoothing in the financial industry, nowcasting of implied volatility means constructing a smooth surface that is consistent with the prices presently observed on a given option market. Option price data arises highly dynamically in ever-changing spatial configurations, which poses a major limitation to foundational machine learning approaches using classical neural networks. While large models in language and image processing deliver breakthrough results on vast corpora of raw data, in financial engineering the generalization from big historical datasets has been hindered by the need for considerable data pre-processing. In particular, implied volatility smoothing has remained an instance-by-instance, hands-on process both for neural network-based and traditional parametric strategies. Our general operator deep smoothing approach, instead, directly maps observed data to smoothed surfaces. We adapt the graph neural operator architecture to do so with high accuracy on ten years of raw intraday S&P 500 options data, using a single model instance. The trained operator adheres to critical no-arbitrage constraints and is robust with respect to subsampling of inputs (occurring in practice in the context of outlier removal). We provide extensive historical benchmarks and showcase the generalization capability of our approach in a comparison with classical neural networks and SVI, an industry standard parametrization for implied volatility. The operator deep smoothing approach thus opens up the use of neural networks on large historical datasets in financial engineering.
@@ 1 Introduction
Options trading experienced phenomenal growth in recent years. In its 2023 trading volume report (Cboe Global Markets, Inc.,, 2024), the CBOE announced the fourth consecutive year of record-breaking volumes on its options exchanges, citing a record-breaking number of transactions for European options on the S&P 500 index. European options are financial derivative contracts that give their holder the right, but not the obligation, to either buy or sell an underlying asset at a predetermined price (the strike) at a predetermined time (the expiry). An option specifying the right to buy (respectively to sell) is called a Call (respectively Put) option. Options are traded on a wide range of underlyings, including stocks, indices, currencies and commodities, and can be used to hedge against or speculate on the price movements of the underlying asset.
@@
A key concept in options trading is the so-called implied volatility, which transforms the nominal price of an option into a conceptually and numerically convenient metric. The implied volatility surface is the collection of implied volatilities as observed at a specific point in time, visualized in three-dimensional space as a surface over the (strike, expiry)-domain. It provides an intuitive representation of the current state of the options market and is crucial for hedging and risk management. The extraction of a smooth surface from quoted option prices is called implied volatility smoothing and allows to infer (or nowcast) theoretical option prices for interpolated strike values and expiry times. It remains one of the key challenges in options trading.
@@
Conventionally, implied volatility smoothing relies on parametric surfaces whose parameters are optimized based on the distance to observed prices while adhering to absence-of-arbitrage conditions, which ensure the consistency of prices extrapolated from the smoothed surface. The development of such ad-hoc models for implied volatility traces back to SVI (Gatheral,, 2004), which models implied volatility slice-wise for each maturity and successfully captures its key features on Equity indices. A continuous interpolation scheme for SVI slices was provided in Gatheral and Jacquier, (2014), yielding a full surface. Nowadays, sophisticated market makers employ custom parametrizations, which can be considered proprietary trading secrets and reduce SVI to a benchmark role.
@@
Regardless of the particular parametric surface model used, the conventional smoothing approach boils down to the continued execution of a numerical optimization routine: A smoothed surface expires as soon as quotes are updated (whenever markets move), necessitating the re-calibration of parameters. Success and duration of this routine is sensitive to initial conditions, search heuristics, and termination criteria, which exposes practitioners to considerable process uncertainties during trading hours (or online). In response, we introduce a novel operator deep smoothing approach, replacing the instance-by-instance optimization with a single evaluation of a neural network. This greatly simplifies online calibration, at the upfront cost of training the network offline from historical data (in the spirit of Hernandez, (2016); Horvath et al., (2021); Liu et al., (2019)). Our unique use of neural operators (Kovachki et al.,, 2023) is fundamentally directed by the observation that the raw inputs for volatility smoothing (the collections of observed volatilities) over time vary in size and spatial arrangement: Options expire, new maturities and strikes become available, and the coordinates of existing options evolve continuously in the domain of the implied volatility surface (Figure 1(b)). This setting excludes classical neural networks – which required fixed-size inputs – from direct application. Neural operators, instead, conceptualize observed data as point-wise discretizations of latent functions in implicit infinite-dimensional function spaces and are well suited for the task.
@@
We introduce operator deep smoothing, a general approach for discretization-invariant data interpolation based on neural operators, and apply it to implied volatility smoothing. Our technique transcends traditional parametric smoothing and directly maps observed volatilities to smoothed surfaces. Comparable neural network based approaches are limited to certain option markets (e.g. FX markets, where options by default spread out on fixed rectilinear grids, as relied upon by Bergeron et al., (2021) for its VAE approach) or require data pre-processing (as in Cont and Vuletić, (2023), which achieves fixed rectilinear grids by linear interpolation of market values, setting aside questions related to no-arbitrage constraints). Instead, our technique novelly adapts the graph neural operator (GNO) architecture (Anandkumar et al.,, 2020) to consistently smooth input data of any size and spatial arrangement. While neural operators have successfully been used in Physics to numerically solve partial differential equations, our application is the first in financial engineering and highlights the values of their discretization-invariance properties, so far rather under-explored. We employ our method on ten years of intraday S&P 500 options data, smoothing more than 60 million volatility datapoints using a single model instance with around 100 thousand trained parameters. We report errors that substantially improve on the SVI industry benchmark and are highly competitive with Ackerer et al., (2020), which performs smoothing by training one classical neural network per volatility surface. We proceed to successfully demonstrate the generalization capabilities of our model for end-of-day options data of the S&P 500 as well as three further major US indices. No data from these three indices has been used for training.
We explore the technical implications of our method in Section 5. Here, we discuss the broader impact of our contributions:
  • Operator Deep Smoothing for Implied Volatility – Our method massively simplifies volatility smoothing, an area where effective methods mean competitive advantage and frequently remain trade secrets. Therefore, we believe that our approach lowers the entry-barrier for effective volatility smoothing even among industry professionals. Practitioners and researchers that are not directly involved in options trading frequently employ rudimentary methods (SVI or linear/spline interpolation). Here, our trained operator, served as a hands-off tool, could provide “cheap” and accurate surfaces for use in downstream tasks. Ultimately, our method may be useful for all participants of option markets. This includes the general public, whose trading in such markets has been increasing substantially (Doherty et al.,, 2023) and which benefits from broadly accessible investment tools.
@@
  • Neural Operators for Discretization-Invariant Interpolation – Our operator deep smoothing approach constitutes the first application of neural operators for interpolation/extrapolation tasks and paves the way to future research on the versatility of the discretization-invariance of neural operators in industrial applications characterized by dynamic and spatially irregular data. At least in financial engineering, surfaces similar to the implied volatility surface (and higher-dimensional equivalents, as for example the volatility cube) are ubiquitous. We expect our technique to be transferable and to streamline and robustify engineers’ and researchers’ algorithms and data pipelines.
@@
The aforementioned SVI was developed for internal use at Merrill Lynch in 1999 and later advocated in Gatheral, (2004). Its extension to surface-based SSVI in Gatheral and Jacquier, (2014) has been eagerly adopted by practitioners, which have since contributed to its robust calibration and generalizations (Corbetta et al.,, 2019; Hendriks and Martini,, 2017; Guo et al.,, 2016). It was augmented in Ackerer et al., (2020) by a multiplicative neural network corrector, based on guided network training by means of no-arbitrage soft constraints from Zheng, (2018). The absence-of-arbitrage conditions for implied volatility surfaces – providing safeguards for option pricing – were formulated in Roper, (2010), and we provide an equivalent formulation, based on Fukasawa, (2012); Lucic, (2021), for practical purposes. In Chataigner et al., (2020) static arbitrage constraints were used to perform option calibration (with an additional regularization technique), which can be considered to be instance-by-instance smoothing of nominal price data. In Bergeron et al., (2021) a classical VAE (variational autoencoder) was applied to implied volatility smoothing on FX markets, where strikes of quoted options are tied to a fixed grid of deltas. [^1] Delta is the derivative of the price with respect to the underlying asset and is standard in FX strike quoting. This specificity of FX markets allows the use of a conventional feedforward neural network based decoder. Recent option calibration approaches based on neural networks have been proposed in Baschetti et al., (2024); Hernandez, (2016); Horvath et al., (2021); Van Mieghem et al., (2023).
A comprehensive account on neural operators is given in Kovachki et al., (2023), unifying previous research on different neural operator architectures and techniques (Anandkumar et al.,, 2020; Li et al.,, 2020). Subsequent developments investigating the expressivity of these architectures as well as their generalizations include Hao et al., (2023); Huang et al., (2024); Lanthaler et al., (2023); Li et al., (2021); Lingsch et al., (2023); Tran et al., (2021).
@@
We review financial concepts and the challenges of implied volatility smoothing in Section 2. In Section 3, we provide a review of neural operators (Section 3.1) and introduce our operator deep smoothing approach for general interpolation tasks (Section 3.2). In Section 4, we perform experiments for implied volatility smoothing of S&P 500 options data. Finally, Section 5 gathers limitations as well as outlooks regarding the use of neural operators for interpolation purposes.
@@
We make code for the paper available at the location https://github.com/rwicl/operator-deep-smoothing-for-implied-volatility. In particular, the code repository contains a general PyTorch (Paszke et al.,, 2019) implementation of the graph neural operator architecture for operator deep smoothing.
@@ 2 Background: Implied Volatility
We consider a market of European options written on an underlying asset, which we observe at a given instant $T_{0}$, and denote the time-$T$ forward price of the underlying asset by $F_{T_{0},T}$.
@@
While not able (any longer) to fit market data, the mathematical tractability of the Black-Scholes model gave rise to the concept of implied volatility: Given a Call option with price $C(T,K)$, its implied volatility $v(\tau,k)$ is defined by $C(T,K)=F_{T_{0},T}\operatorname{BS}(\tau,k,v(\tau,k))$. By using time-to-expiry/log-moneyness coordinates, implied volatility provides a universal way to consistently compare the relative expensiveness of options of across different strikes, expiries, underlyings, and interest rate environments. Its characteristic shape helps traders make intuitive sense of the instantaneous state of option markets relative to a (flat) Black-Scholes model baseline.
@@
This refers to fitting a smooth surface $\hat{v}\colon(0,\infty)\times\mathbb{R}\to(0,\infty)$ to a collection $\mathbf{v}=\{v(\tau_{l},k_{l})\}_{l=1}^{p}$ of observed implied volatilities. Naive strategies such as cubic interpolation are ill-fated: Surfaces generated by such interpolation rules will in general correspond to Call option prices that are exploitable by so-called arbitrage, namely cost-less trading strategies generating a guaranteed profit. In option markets, an arbitrage is called static when set up solely from fixed positions in options and a dynamically (but a finite number of times) readjusted position in the underlying. Beyond simple interpolations, practitioners have devised ad-hoc parametrizations for implied volatility, in particular the aforementioned SVI, which are not expected to perfectly match all reference prices. Instead, the model parameters are optimized with respect to an objective function that measures market price discrepancy and includes penalization terms ruling out static arbitrage. These penalization terms are commonly formulated on the basis of the following theorem, which summarizes the shape constraints of the implied volatility surface (Gatheral and Jacquier,, 2014; Lucic,, 2021; Roper,, 2010).
@@
Condition 2.1(ii) is equivalent to prices increasing in maturity (uncertain increases as time passes), while Condition 2.1(i) arises when computing the implied probability density $f_{\tau}$ of the underlying:
@@
Since a density needs to be non-negative, (3) explains why Condition 2.1(i) above is required.
(a) Relative trading volume per quoted interval, averaged over S&P 500 dataset 2012-2021. Left: Rectangular domain w.r.t.  time-to-expiry/log-moneyness (99.9% of trading volume with maturities ¡1 year). Right: Rectangular domain w.r.t. transformed coordinates (96.6% of trading volume while more evenly populated); boundary delineated in Left.
@@ 3.1 Background: Neural Operators
An input-output relationship of data $\mathbf{a}\mapsto\mathbf{u}$ is then ”really” described by an operator $F\colon\mathcal{A}\to\mathcal{U}$ between function spaces $\mathcal{A}$ and $\mathcal{U}$. Neural operators are abstract neural network architectures $F^{\theta}\colon\mathcal{A}\to\mathcal{U}$, with implementations that integrate (4) consistently across the variable discretization $\pi$.
@@
We review the core concepts of neural operators from Kovachki et al., (2023). Let $D$ be a bounded domain in $\mathbb{R}^{d}$ and $\mathcal{A}$ and $\mathcal{U}$ be Banach spaces of functions mapping from $D$ to $\mathbb{R}^{c_{\text{in}}}$ and $\mathbb{R}^{c_{\text{out}}}$, respectively. Neural operators are finitely parametrized mappings $F^{\theta}\colon\mathcal{A}\to\mathcal{U}$ with universality for continuous target operators and with discretization-invariant implementations $\tilde{F}^{\theta}$. In the space $C(\mathcal{A},\mathcal{U})$ topologized by uniform convergence on compacts, the architecture $F^{\theta}$ is called universal if $\{F^{\theta}\}_{\theta\in\Theta}$ is dense in $C(\mathcal{A},\mathcal{U})$, with $\Theta$ the parameter set. An implementation of $F^{\theta}$ is an algorithm $\tilde{F}^{\theta}$ which accepts observed data $\mathbf{a}=a|_{\pi}$ and outputs a function $u\in\mathcal{U}$ and is such that $\tilde{F}^{\theta}_{\pi}(\makebox[4.30554pt]{{$\cdot$}})=\tilde{F}^{\theta}(\makebox[4.30554pt]{{$\cdot$}}|_{\pi})\in C(\mathcal{A},\mathcal{U})$. Now, $\tilde{F}^{\theta}$ is called discretization-invariant if $\lim_{n\uparrow\infty}\tilde{F}^{\theta}_{\pi^{(n)}}=F^{\theta}$ in $C(\mathcal{A},\mathcal{U})$, given a discrete refinement [^3] A discrete refinement of $D$ is a nested sequence $(\pi^{(n)})_{n\in\mathbb{N}}$ of discretizations of $D$ for which for every $\varepsilon>0$, there exists $N\in\mathbb{N}$ such that $\{\mathbb{B}_{\mathbb{R}^{d}}(x,\varepsilon)\colon x\in\pi^{(N)}\}$ covers $D$. of $D$.